# Universal projection theorems with applications to multifractal analysis and the dimension of every ergodic measure on self-conformal sets simultaneously

### v2

@@ Abstract.
We prove a universal projection theorem, which gives conditions on a parametrized family of maps $\Pi_{\lambda}:X\to\mathbb{R}^{d}$ and a collection ${\mathcal{M}}$ of measures on $X$ under which for almost every $\lambda$ equality $\dim_{H}\Pi_{\lambda}\mu=\min\{d,\dim_{H}\mu\}$ holds for all measures $\mu\in{\mathcal{M}}$ simultaneously (i.e. on a full measure set of $\lambda$’s independent of $\mu$). We require family $\Pi_{\lambda}$ to satisfy a transversality condition and collection ${\mathcal{M}}$ to satisfy a new condition called relative dimension separability. Under the same assumptions, we also prove that if the Assouad dimension of $X$ is smaller than $d$, then for almost every $\lambda$, projection $\Pi_{\lambda}$ is nearly bi-Lipschitz (i.e. with pointwise $\alpha$-Hölder inverse for every $\alpha\in(0,1)$) at $\mu$-a.e. $x$, for all measures $\mu\in{\mathcal{M}}$ simultaneously. Our setting is general enough to include families of orthogonal projections, natural projections corresponding to conformal iterated functions systems, as well as non-autonomous or random IFS.
As applications, we study families of ergodic measures on self-conformal sets. We prove that given a parametrized family of contracting conformal $C^{1+\theta}$ IFS ${\mathcal{F}}_{\lambda}=\{f_{i}^{\lambda}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}\}_{i\in{\mathcal{A}}}$ satisfying the transversality condition, for almost every parameter $\lambda$ one has $\dim_{H}\Pi_{\lambda}\mu=\min\{d,\frac{h(\mu)}{\chi(\lambda,\mu)}\}$ for all ergodic shift-invariant measures on ${\mathcal{A}}^{\mathbb{N}}$ simultaneously, i.e. with a common exceptional set of parameters $\lambda$ of zero measure (here $h(\mu)$ is the entropy of $\mu$ and $\chi(\lambda,\mu)$ is its Lyapunov exponent).
As a second application, we prove that given a self-similar system $\{x\mapsto\lambda_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ on the line with similarity dimension $s_{0}<1$, Lebesgue almost every choice of translations $t_{i}$ has the property that the multifractal formalism holds on the full spectrum interval $\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}\right]$ for every self-similar measure corresponding to a probability vector $(p_{i})_{i\in{\mathcal{A}}}$ simultaneously.
We also prove that the dimension part of the Marstrand-Mattila projection theorem holds simultaneously for the collection of all ergodic measures on a self-conformal set satisfying the strong separation condition and for the collection of all Gibbs measures on a self-conformal set (without any separation).
@@
B. Bárány acknowledges support from the grant NKFI FK134251. B. Bárány and K. Simon was supported by the grants NKFI K142169 and KKP144059 "Fractal geometry and applications". A. Śpiewak was partially supported by the National Science Centre (Poland) grant 2020/39/B/ST1/02329. We are grateful to Thomas Jordan and Boris Solomyak for useful discussions which have inspired this work.
@@ 1. Introduction
(1.1) $\overline{\dim}_{H}(P_{V}\mu)=\min\{d,\overline{\dim}_{H}\mu\}\text{ and }\underline{\dim}_{H}(P_{V}\mu)=\min\{d,\underline{\dim}_{H}\mu\}\text{ for }\gamma\text{-a.e. }V\in\operatorname{Gr}(d,n),$
@@
where $\overline{\dim}_{H}$ and $\underline{\dim}_{H}$ denote the upper and lower Hausdorff dimensions (see Section 4 for definitions), $\operatorname{Gr}(d,n)$ is the Grassmannian manifold of $d$-dimensional linear subspaces in $\mathbb{R}^{n}$ endowed with the unique $O(n)$-invariant probability measure $\gamma$ (see [Mat95, Section 3.9]) and $P_{V}$ denotes the orthogonal projection onto $V\in\operatorname{Gr}(d,n)$. These works have been the "precursor" of the transversality method to handling the overlaps for iterated function systems.
Let ${\mathcal{A}}$ be a finite collection of indices and let =${\mathcal{F}}=\{f_{i}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}\}_{i\in{\mathcal{A}}}$ be a finite collection of contracting maps, called iterated function system (IFS), and let $p=(p_{i})_{i\in{\mathcal{A}}}$ be a probability measure on ${\mathcal{A}}$. Then there exists a unique non-empty compact set $\Lambda$ and a unique compactly supported probability measure $\nu$ such that $\Lambda$ is invariant with respect to ${\mathcal{F}}$ in the sense that $\Lambda=\bigcup_{i\in{\mathcal{A}}}f_{i}(\Lambda)$ and $\nu$ is stationary, i.e. $\nu=\sum_{i\in{\mathcal{A}}}p_{i}(f_{i})_{*}\nu$ see Hutchinson [Hut81]. There is a natural correspondence between the points of $\Lambda$ and infinite sequences formed by ${\mathcal{A}}$. More precisely, there is a Lipschitz map $\Pi\colon{\mathcal{A}}^{\mathbb{N}}\to\Lambda$ defined by $\Pi(u_{1},u_{2},\ldots)=\lim\limits_{n\to\infty}f_{u_{1}}\circ\cdots\circ f_{u_{n}}(0)$, called the natural projection. It satisfies $\Lambda=\Pi({\mathcal{A}}^{\mathbb{N}})$ and stationary measure $\nu$ can be obtained as the projection of the Bernoulli measure $p^{\mathbb{N}}$ on ${\mathcal{A}}^{\mathbb{N}}$, i.e. $\nu=\Pi(p^{{\mathbb{N}}})$. Calculating the dimension of $\Lambda$ and $\nu$ is a challenging problem in general, even in the case, when the maps of ${\mathcal{F}}$ are similarities or conformal maps. In case of a separation condition, for example if the open set condition holds, then the dimension (both box and Hausdorff) of $\Lambda$ equals to the similarity dimension, and also, the dimension of $\nu$ is given by the ratio of the entropy and the Lyapunov exponents - see Example 3.3 for details and precise definitions.
The transversality method was first applied by Pollicott and Simon [PS95] to the case of iterated function systems, and later several generalizations have appeared, see for instance Peres and Solomyak [PS96], Solomyak [Sol98], Peres and Schlag [PS00], Peres, Schlag and Lindenstrauss [LPS02], Simon, Solomyak and Urbański [SSU01a, SSU01b], Bárány, Simon, Solomyak and Śpiewak [BSSŚ22, BSSŚ24]. See also [Sol23] for a recent survey and [BSS23] for an in-depth discussion. Roughly speaking, the dimensional parts of these statements can be summarized as follows. Let ${\mathcal{F}}_{\lambda}=\{f_{i}^{\lambda}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}\}_{i\in{\mathcal{A}}},\ \lambda\in U$ be a family of parametrized conformal iterated function systems, which satisfies the transversality condition with respect to a probability measure $\eta$ on $U$ (see precise definition later). Let $\Pi_{\lambda}$ be the natural projection map corresponding to ${\mathcal{F}}_{\lambda}$ and let $\mu$ be an ergodic left-shift invariant measure ${\mathcal{A}}^{\mathbb{N}}$. Then
(1.2) $\dim_{H}(\Pi_{\lambda}\mu)=\min\left\{d,\frac{h(\mu)}{\chi(\mu,\lambda)}\right\}\text{ for }\eta\text{-a.e. }\lambda$
where $h(\mu)$ denotes the entropy of $\mu$ and $\chi(\mu,\lambda)$ denotes the Lyapunov exponent of $\mu$ with respect to ${\mathcal{F}}_{\lambda}$. As explained in Example 3.3, any result of the type (1.2) can be seen as an IFS analogue of the Marstrand-Mattila projection theorem (1.1), by endowing ${\mathcal{A}}^{\mathbb{N}}$ with a proper metric.
@@
The goal of this paper is to study conditions under which (1.1) and (1.2) can be proved to hold for all measures in a given class of measures ${\mathcal{M}}$ simultaneously, i.e. with a measure zero set of exceptional projections independent of measure $\mu\in{\mathcal{M}}$. For instance, in the context of orthogonal projections we look for assumptions on a set ${\mathcal{M}}\subset{\mathcal{M}}_{{\mathrm{fin}}}(\mathbb{R}^{n})$ of finite Borel measures on $\mathbb{R}^{n}$ so that (1.1) can be improved to

### v3

@@ Abstract.
We prove a universal projection theorem, giving conditions on a parametrized family of maps $\Pi_{\lambda}:X\to\mathbb{R}^{d}$ and a collection $\mathcal{M}$} of measures on $X$ under which for almost every $\lambda$ equality $\dim_{H}\Pi_{\lambda}\mu=\min\{d,\dim_{H}\mu\}$ holds for all measures $\mu\in{\mathcal{M}}$ simultaneously (i.e. on a full measure set of $\lambda$’s independent of $\mu$). We require family $\Pi_{\lambda}$ to satisfy a transversality condition and collection $\mathcal{M}$} to satisfy a new condition called relative dimension separability. Under the same assumptions, we also prove that if the Assouad dimension of $X$ is smaller than $d$, then for almost every $\lambda$, projection $\Pi_{\lambda}$ is nearly bi-Lipschitz (i.e. with pointwise $\alpha$-Hölder inverse for every $\alpha\in(0,1)$) at $\mu$-a.e. $x$, for all $\mu\in{\mathcal{M}}$ simultaneously. Our setting encompasses families of orthogonal projections, natural projections corresponding to conformal iterated function systems, and non-autonomous or random IFS.
As applications, we provide novel results on the multifractal analysis, giving formula for the Hausdorff dimension of a level set of the local dimension for a typical (w.r.t the translation parameter) self-similar measure on the line, valid for the full range spectrum (including the decreasing part of the spectrum; previous results were covering only the increasing part).
Among another applications, we prove that given a parametrized contracting conformal IFS satisfying the transversality condition, for almost every parameter the dimension formula holds for all ergodic shift-invariant measures simultaneously. We also prove that the dimension part of the Marstrand’s projection theorem holds simultaneously for the collection of all ergodic measures on a strongly separated self-conformal set and for the collection of all Gibbs measures on a self-conformal set (without any separation).
@@
B. Bárány acknowledges support from the grant NKFI FK134251. B. Bárány and K. Simon were supported by the grants NKFI K142169 and KKP144059 "Fractal geometry and applications". A. Śpiewak was partially supported by the National Science Centre (Poland) grant 2020/39/B/ST1/02329. We are grateful to Thomas Jordan and Boris Solomyak for useful discussions which have inspired this work.
@@ 1.1. Marstrand’s projection theorem
(1.1) $\dim_{H}(P_{V}\mu)=\min\{d,\dim_{H}\mu\}\text{ for }\gamma\text{-a.e. }V\in\operatorname{Gr}(d,n),$
@@
where $\dim_{H}$ denotes the Hausdorff dimension111in fact, for general measures one should consider the upper and lower Hausdorff dimensions separately. For simplicity, we omit this subtlety in the introduction., $\operatorname{Gr}(d,n)$ is the Grassmannian manifold of $d$-dimensional linear subspaces in $\mathbb{R}^{n}$ endowed with the unique $O(n)$-invariant probability measure $\gamma$ (see [Mat95, Section 3.9]) and $P_{V}$ denotes the orthogonal projection onto $V\in\operatorname{Gr}(d,n)$ - see Section 4 for definitions.
Marstrand’s projection theorem has been generalized and extended in several directions, with the methods behind its proof becoming an underlying paradigm for many areas of research in geometric measure theory [Mat95, Mat15, FFJ15], fractal geometry [PS95, Sol95, Sol23] or dynamical systems [SYC91, Tsu15]. The method is often referred to as the transversality technique. Due to its wide range of applications, it has been extended to generalized projection schemes, which apply to several settings at once, see e.g. [Sol98, PS00, LPS02, BT23, BSS23]. The aim of this paper is to improve the strength of this technique in the setting of generalized projections and utilize it for novel applications in geometric measure theory and fractal geometry.
1.2. Main results - universal projection theorem
The main result of this paper is Theorem 2.7. Let us explain it first in a non-technical manner. Let $\Pi_{\lambda}:X\to\mathbb{R}^{d}$ be a collection of Lipschitz maps (projections) parametrized by a parameter $\lambda\in U$ and satisfying the transversality condition with respect to a measure $\eta$ on $U$ (see Section 2 for precise definitions and assumptions). A direct extension of (1.1) to such a general setting states that given a finite measure $\mu$ on $X$ the following holds
(1.2) $\dim_{H}\Pi_{\lambda}\mu=\min\{d,\dim_{H}\mu\}\text{ for }\eta\text{-a.e. }\lambda\in U,$
@@ 1.2. Main results - universal projection theorem
see e.g. [BSS23, Theorem 6.6.2]. The first goal of this paper is to study conditions under which (1.1), and more generally (1.2), hold for all measures in a given class of measures $\mathcal{M}$} simultaneously, i.e. with a measure zero set of exceptional projections independent of measure $\mu\in{\mathcal{M}}$. More precisely, we aim at finding conditions on $\mathcal{M}$} guaranteeing that
(1.3) $\eta\left(\left\{\lambda\in U:\underset{\mu\in{\mathcal{M}}}{\exists}\dim_{H}\Pi_{\lambda}\mu\neq\min\{d,\dim_{H}\mu\}\right\}\right)=0$
holds. Clearly, this question is relevant only for uncountable families $\mathcal{M}$} and it is easy to find examples of collections $\mathcal{M}$} for which (1.3) does not hold (see Example 3.2 for the case of orthogonal projections). Our main result - Theorem 2.7 - identifies a condition on $\mathcal{M}$}, which we call relative dimension separability (see Definition 2.3), guaranteeing that (1.3) holds. We refer to the result as the universal projection theorem222the name universal projection theorem is inspired by seminal results and constructions known in information theory as universal source codings, providing compression algorithms operating in an optimal rate for any source distribution in a given class, without any prior knowledge of this distribution, see e.g. [CT06, Chapter 13]. A famous example is the Lempel-Ziv coding [ZL77, ZL78] achieving optimal rate for any stationary ergodic source distribution [CT06, Theorem 13.5.3].. We shall emphasize, that the projection scheme studied in this paper does not require any extra assumptions than the classical ones, hence the simultaneous result (1.3) holds essentially whenever the standard transversality technique can be applied.
We also study the same problem in the context of embeddings, corresponding to the injectivity properties of the projections. This is an important topic in analysis and geometric measure theory, studied by numerous authors [Mn81, SYC91, EFNT94, BAEFN93, HK99, Ols02, RS19]. In recent years there has been an increasing interest in almost surely injective embeddings of measures, due to its applications to compression of analog signals [WV10, ABDL+19, GŚ20] and time-delayed embeddings of dynamical systems [SSOY98, BGŚ20, BGŚ22]. Of particular importance is establishing guarantees on the regularity of the embedding [HK99, Ols02, BGŚ23, Śpi25]. In the fractal setting, strongest regularity properties have been obtained in terms of the Assouad dimension [Ols02] (see also [Rob11]), guaranteeing that the inverse to the orthogonal projection of a compact set $X\subset\mathbb{R}^{N}$ onto a typical $k$-dimensional plane is almost bi-Lipschitz (in particular: $\alpha$-Hölder for every $\alpha\in(0,1)$) provided that $\dim_{A}(X-X)<k$ (here $\dim_{A}$ denotes the Assouad dimension, see Definition 4.2). It was shown in [BGŚ23, Theorem 1.7.(iii)], that given a probability measure $\mu$ on $X$, a weaker condition $\dim_{A}X<k$ suffices to guarantee that the inverse is defined $\mu$-almost everywhere and it is pointwise almost bi-Lipschitz at $\mu$-a.e. $x\in X$. In Theorem 2.7 we improve this result to a universal version for general projection schemes. In a spirit similar to (1.3), we prove in Theorem 2.7 that the following holds for relative dimension separable collections of measures $\mathcal{M}$} on $X$
(1.4) $\text{if }\dim_{A}X<d,\text{ then }\eta\left(\left\{\lambda\in U:\underset{\mu\in{\mathcal{M}}}{\exists}\ \Pi_{\lambda}\text{ is not }\mu\text{-nearly bi-Lipschitz}\right\}\right)=0,$
with the precise definition of the $\mu$-nearly bi-Lipschitz property given in Definition 2.6. Establishing the universal nearly bi-Lipschitz property (1.4) is a crucial step for further applications. As we explain later, in the context of iterated function systems, the nearly bi-Lipschitz property is equivalent to a new and non-trivial separation condition called exponential distance from the enemy (EDE) (see Definition 3.9 and Proposition 3.10). Its usefulness is verified by novel applications to the multifractal analysis, elaborated on below. See also Remark 3.11.
1.3. Applications I - multifractal analysis of self-similar measures
Our most important application of the universal projection theorem described above is to the multifractal analysis of self-similar measures. To explain it, let us first introduce the general setting of iterated functions systems.
Let $\mathcal{A}$} be a finite collection of indices and let ${\mathcal{F}}=\{f_{i}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}\}_{i\in{\mathcal{A}}}$ be a finite collection of (strictly) contracting maps, called iterated function system (IFS), and let $p=(p_{i})_{i\in{\mathcal{A}}}$ be a probability measure on $\mathcal{A}$}. Then there exists a unique non-empty compact set $\Lambda$ and a unique compactly supported probability measure $\nu$ such that $\Lambda$ is invariant with respect to $\mathcal{F}$} in the sense that $\Lambda=\bigcup\limits_{i\in{\mathcal{A}}}f_{i}(\Lambda)$ and $\nu$ is stationary, i.e. $\nu=\sum_{i\in{\mathcal{A}}}p_{i}(f_{i})_{*}\nu$ - see Hutchinson [Hut81]. Measure $\nu$ is called self-similar if $\mathcal{F}$} consists of similarity maps. The multifractal analysis studies the level-set function $\alpha\mapsto\dim_{H}\left(\left\{x:d(\nu,x)=\alpha\right\}\right)$. It can be seen as a fine-scale analysis of the geometry of measure $\mu$ and originates from the study of the distribution of singularities on strange attractors in a chaotic, nonlinear dynamics [HJK+86]. Later, a rigorous mathematical approach has been developed and became one of the central topics in fractal geometry, see e.g. [Pes98] for a more detailed discussion. A basic heuristics for a self-similar measure corresponding to a probability vector $p$ and an IFS ${\mathcal{F}}=\{f_{i}(x)=\lambda_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ is that one expects the formula
(1.5) $\dim_{H}\left(\left\{x:d(\nu,x)=\alpha\right\}\right)=\inf_{q\in\mathbb{R}}(\alpha q+T(q)),$
to hold, where for $q\in\mathbb{R}$ we define $T(q)$ to be the unique real number satisfying $\sum\limits_{i\in{\mathcal{A}}}p_{i}^{q}|\lambda_{i}|^{T(q)}=1$ (the quantity $T(q)$ is related to the $L^{q}$-spectrum of the self-similar measure, see Remark 3.16). The maximal range of $\alpha$’s on which one can expect (1.5) to hold is $\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}\right]$. Arbeiter and Patschke [AP96] showed that (1.5) holds for the full range of $\alpha$’s if $\mathcal{F}$} satisfies the Open Set Condition. With the aid of the universal projection theorem we can prove the following result, addressing the overlapping case for typical parameters.
Theorem 1.1.
Fix $\lambda_{i}\in(-1,1)\setminus\{0\},i\in{\mathcal{A}}$ for a finite set $\mathcal{A}$}. For each $t=(t_{i})_{i\in{\mathcal{A}}}\in\mathbb{R}^{{\mathcal{A}}}$ and a probability vector $p=(p_{i})_{i\in{\mathcal{A}}}$, let $\nu_{t,p}$ be the self-similar measure corresponding to the IFS ${\mathcal{F}}_{t}=\{f_{i}(x)=\lambda_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ on $\mathbb{R}$ and the probability vector $p$. The following holds for Lebesgue almost every $t\in\mathbb{R}^{\mathcal{A}}$. If the similarity dimension $s_{0}=s({\mathcal{F}}_{t})$ of ${\mathcal{F}}_{t}$ satisfies $s_{0}<1$, then equality
$\dim_{H}\left(\left\{x:d(\nu_{t,p},x)=\alpha\right\}\right)=\inf_{q\in\mathbb{R}}(\alpha q+T(q))$
holds for every $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}\right]$ and every probability vector $p$ such that $p_{i}\neq|\lambda_{i}|^{s_{0}}$ for some $i\in{\mathcal{A}}$.
See also Theorem 3.15 for the multidimensional version. Let us explain the significance of the above result. So far, the strongest result on the multifractal analysis of self-similar measures with overlaps is the one by Barral and Feng [BF21, Theorem 1.2, Remark 7.3] (improving upon their earlier result [BF13, Theorem 1.3, Theorem 6.4]), which establishes (1.5) for self-similar measures with $s_{0}<1$ and satisfying the Exponential Separation Condition (recall that it is satisfied for $t\in\mathbb{R}^{\mathcal{A}}$ outside of a set of positive Hausdorff codimension, so in particular for almost every $t\in\mathbb{R}^{\mathcal{A}}$), for $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\frac{\sum_{i}|\lambda_{i}|^{s_{0}}\log p_{i}}{\sum_{i}|\lambda_{i}|^{s_{0}}\log|\lambda_{i}|}\right]$. This range corresponds to the increasing part of the spectrum. The remaining range $\left[\frac{\sum_{i}|\lambda_{i}|^{s_{0}}\log p_{i}}{\sum_{i}|\lambda_{i}|^{s_{0}}\log|\lambda_{i}|},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}\right]$ (the decreasing part of the spectrum), uncovered by their results, corresponds to $\alpha$’s for which the infimum in (1.5) is attained for $q<0$ and hence is related to studying the $L^{q}$-spectrum of the self-similar measure for negative $q$. This is known to be notoriously difficult and there are no corresponding results available for typical self-similar measures with overlaps in this range. With the use of the nearly bi-Lipschitz property from Theorem 2.7 we are able to circumvent this problem. According to our best knowledge, Theorem 1.1 is the first result on the multifractal spectrum of typical self-similar measures with overlaps, which covers the full spectrum. Let us emphasize that while the introduction of additive combinatorics methods of Hochman [Hoc14] and Shmerkin [Shm19] to the study of self-similar measures have led to deep breakthroughs in recent years (including results of Barral and Feng [BF21] on multifractal formalism for self-similar measures), these methods have not been yet able to provide results on the multifractal analysis in the full range $\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}\right]$. The progress in Theorem 1.1 is based on the classical transversality technique, with establishing (1.4) as the crucial ingredient. It is easy to extend our results to the Birkhoff spectra of continuous potentials.
1.4. Applications II - dimension of every ergodic measure on self-conformal sets simultaneously
The transversality method was first applied to the case of iterated function systems by Pollicott and Simon [PS95], and later several generalizations have appeared, see for instance [PS96, Sol98, PS00, LPS02, SSU01a, SSU01b, BSSŚ22, BSSŚ24]. See also [Sol23] for a recent survey and [BSS23] for an in-depth discussion. Roughly speaking, the dimensional parts of these statements can be summarized as follows. Let ${\mathcal{F}}_{\lambda}=\{f_{i}^{\lambda}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}\}_{i\in{\mathcal{A}}},\ \lambda\in U$ be a family of parametrized conformal ($C^{1+\theta}$) IFS, which satisfies the transversality condition with respect to a probability measure $\eta$ on $U$ (see Example 3.7 for the precise assumptions). Let $\Pi_{\lambda}:{\mathcal{A}}^{\mathbb{N}}\to\mathbb{R}^{d},\ \Pi_{\lambda}(\omega)=\lim\limits_{n\to\infty}f_{\omega_{1}}\circ\cdots\circ f_{\omega_{n}}(x)$ be the natural projection map corresponding to ${\mathcal{F}}_{\lambda}$ and let $\mu$ be an ergodic left-shift invariant measure on $\mathcal{A}$$^{\mathbb{N}}$. Then (see e.g. [BSS23, Theorem 14.4.2])
(1.6) $\dim_{H}(\Pi_{\lambda}\mu)=\min\left\{d,\frac{h(\mu)}{\chi(\mu,\lambda)}\right\}\text{ for }\eta\text{-a.e. }\lambda,$
where $h(\mu)$ denotes the entropy of $\mu$ and $\chi(\mu,\lambda)$ denotes the Lyapunov exponent of $\mu$ with respect to ${\mathcal{F}}_{\lambda}$. With the use of our methods, we can improve this classical result to a version holding simultaneously for all ergodic measures on $\mathcal{A}$$^{\mathbb{N}}$: